OpenStudy (anonymous):

I don't understand how to go from Step 1 to Step 2. Equations below.

OpenStudy (anonymous):

\[\frac{ dx }{ x(2-x) }=dt\] and this is equivalent to: \[\frac{ dx }{ x } + \frac{ dx }{ 2-x }=2dt\] Please tell me why that is!

OpenStudy (anonymous):

my guess, without computing, is that it is partial fractions

OpenStudy (anonymous):

dis is partial fraction.u will get dis in integration chapter.just read the theory and example part.

OpenStudy (anonymous):

\[\frac{1}{x(2-x)}=\frac{A}{x}+\frac{B}{2-x}\] and my guess is that \(A=B=\frac{1}{2}\)

OpenStudy (anonymous):

in fact i can see it eyeball the answer and you will get it

OpenStudy (anonymous):

Thank you. Is this partial fraction covered in Calculus I or II? I have seen it done before but didnt know that it was called partial fraction.

OpenStudy (anonymous):

in this type of problem u have to devide it in 2 parts.first fraction contains A and second one contains B

OpenStudy (anonymous):

often this is a technique taught in pre-calc, but forgotten until calc 2 when you integrate

OpenStudy (anonymous):

I see. Should I be able to find this in Paul's math notes? Thank you both.

OpenStudy (anonymous):

Found it.